OpenStudy (anonymous):

Is the coordinate (1, 2) a solution of the system below? x + 2y = 5 y = x + 1 Yes No

OpenStudy (♪chibiterasu):

(x, y) = (1, 2) Replace the variables with the corresponding numbers. If they are both equal in the end, that is the solution: 1 + 2(2) = 5 2 = 1 + 1 Think you can do the rest? :3

OpenStudy (anonymous):

is it yes or no

OpenStudy (anonymous):

i think its yes

OpenStudy (♪chibiterasu):

Yep. :3 1 + 2(2) = 5 1 + 4 = 5 5 = 5 2 = 1 + 1 2 = 2

OpenStudy (anonymous):

okay thank you